Ops/Comm Scheduling meeting

Location : CR21 @ 14:00 pm on January 17th

Phone access : ReadyTalk number
US 866-740-1260
UK 0800-496-0576
Swiss 0800-70-53-51
Italy 800182592
Access Code: 2155402

  • CC Magnet status
  • Hydrogen status
  • SMs, autoSMS, ssh-lockout, and iocaccess (Pierrick)
  • Issues arising from previous running
  • Shutdown work in Jan and Feb: present version
  • Operations :
    • Shifts open from Feb 13 to March 6. Signup is slow. Please sign up. Run plans are in development.
    • Organisation of controls system check from cold to data-taking : Jan 23 - 24 (Ajit/Ed)
  • AOB:
    • AutoSMS experts list


  • CC Magnet status (JBoehm)
    All 3 CC magnets are now in flip mode; cold and full of helium. FC was run up to 165 A in flip mode on Thursday. SSD M2 was run up to 50 A. SSU not run (so Diffuser test remains).
    Everything is now being run into shorts for the CC IOC testing.
    The lead heaters have been switched off on SSD - monitoring to see if any effect.
    Ajit will transfer the C&M to run from PRO; PMH will continue to work in DEV instance without affecting anything. Code freeze of PRO this Friday.
    JBoehm has written the link-box wiring documents but need revision - waiting on DL for drawings.
    JBoehm to get detailed step-by-step instructions for SS operation from Pierrick, to put these into operating instructions (similar to existing FC docs).
    There was an extended discussion: it's not clear if the low-level understanding of SS operation is being effectively transferred from Pierrick, or if the right balance between his spending time on low-level documentation vs. finishing the Cooling Channel IOC is being met. This should be covered at MIPO not here.
    The Windows licence missing from the QPS machines has now been installed.
  • Hydrogen status (VBayliss)
    The R9 LH2 system is still cold; trying to source sufficient Neon.
    Have tested the heater and confirmed the heat loads on the cryocooler.
  • Issues arising from previous running (DC)
    Need a comprehensive list of encountered issues, else can't assume they will be addressed before running resumes. We are now well past halfway through the shutdown.
    PHodgson has spoken to Ed and Melissa. Durga is working through eLog regarding Alarm-Handler-related items.
    • Recent compressed air failure: need to "clone" PV's and add to SM spreadsheets for inclusion by next run.
      Note that quantities shared among sub-systems should each have a cloned PV with appropriate limits which is handled in that sub-system's ALH via its spreadsheet; this is so that not-very-trained people can see all systems affected by the loss of a service and contact relevant experts.
      MICE' air compressor (the big blue one) should have cut in when the air failed but appears to have instead turned off soon after - this is under investigation.
    • The reported search drop off the PPS was re-created: removing RF Cage key did not drop the search. Thus, it still remains unknown what caused the original incident.
    • Presently still a lot of ALH activations going off
  • Shutdown work in Jan and Feb ( official version )
    Safety Tour outcomes:
    • Remove pre-PPS chains and locks from doors
    • Remove cryo heads to storage
    • Noise survey needs refreshing
    • (Electrical thing on floor behind DS fridge is the interface unit to the Test Cap for the Linde fridge)

    Last Wed. the conventional magnets were all run together at full power: process water stable just below 22°C (38°C return from cryocompressor) with added 195 kW and no air-blast active.
    This week's jobs:

    • installing pipework to connect loading bay chiller
    • re-plumbing of heat exchanger in trench to counter-flow configuration

    Ajit & Ed control systems cold-start test on 23rd and 24th - Ajit will discuss with Henry and Mike tomorrow
    The PPS functional test will be on Wed. 25th January - MICE Hall closed.
    ODH sensor service will be on 30th January.
    DL down en masse week 30th January for electrical work.
    Diffuser test needs to be re-scheduled for February

  • Operations
    Shifts open from Feb 13 to March 6. Signup is slow.
    Controls system check - see above.
    Expert-led Startup: 10th to 13th February. It was noted that this includes a weekend - not all experts will be available. Hope of getting ISIS machine-physics beam time - this is a negotiation between Steve Boyd and the ISIS Machine Scheduling Meeting.
  • AOB

    Neutron Monitor débâcle: no update from ISIS about ours
    DC needs to ensure that the temporary one is still calibrated, tested, and correct alarm limits in force.

    MICE Safety Management Meeting on 2nd February: this is the PPD-led review; people may be asked for input.

    There is now a MOM C&M panel to assist with re-starts after interventions by sub-system experts, which includes lists of numbers to which SMS alerts are to be sent. The numbers associated with each sub-system need to be confirmed; however the overall way this is to work needs to be discussed MICE-wide.


  • JBoehm to get SS operating info from PHanlet
  • DC to collate list of issues arising from previous running
  • Status of miceraid5 disk cleanup (stands)
  • AKurup to make and circulate plan for C&M cold-start test
  • SBoyd - ISIS machine-physics beam time
  • DC to ensure Health Physics have checked temporary monitor

Updated by Nebrensky, Henry over 7 years ago · 11 revisions